OpenStudy (anonymous):

hello! i would like assistance on how to solve a system of three equations in three unknowns using the echelon method

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

2x+y+z=9 -x-y+z=1 3x-y+z=9